Update for the Week Ending March 11, 2016

Site Remediation Continues

On-going activities:

  • Continued the application of water and calcium chloride for dust control;
  • Continued the application of odor-suppressant foam and BioSolve for dust and odor control;
  • Maintained erosion and sediment (E&S) controls (Installed waddles - wood chip filled fabric tubes - around the Site perimeter to prevent silt and sediment from leaving the Site and covered all stockpiles);
  • Continued operation of the WTP;
  • Continued pre-trenching and demolition of subsurface obstructions in the alignment of the Support of Excavation (SOE) wall;
  • Continued the lining, loading, and shipment of impacted soils to Bayshore Soil Management for treatment and disposal:
  • Initiated full scale Bucket Soil Mixing (BSM) In-Situ Stabilization (ISS) in the Holder No. 2 tank; and
  • Mobilized a drill rig and initiated pre-drilling for deep obstruction identification and removal along the alignment of the SOE wall.

Recently completed activities:

  • Completed the ISS BSM stabilization of five of twenty cells (six of twenty completed to date) in the Holder No. 2 tank; and
  • Lined, loaded, and shipped trucks of impacted soils to Bayshore Soil Management and Clean Earth of North Jersey for treatment and disposal:
    • 17 trucks; 445.92 tons shipped last week; and
    • 1,265.15 tons shipped to date.

Upcoming activities:

  • Ship removed asphalt for recycling;
  • Crush clean, reusable concrete for use as backfill in the deep excavation; and
  • Install dewatering wells and initiate deep ground water extraction.

Community Air Monitoring Plan Summary

Community air monitoring, initiated on Tuesday, December 1, 2015, continues as intrusive (ground disturbance) activities are conducted.

There were no fifteen-minute Action Limit exceedances of Total Volatile Organics (TVOCs) or odors during this reporting period.

There was one instance of the fifteen-minute average concentrations of Particulate Matter (PM-10) [dust] exceeding the Action Limit on Friday, March 4, 2016, during the delivery of clean slag/cement to the grout plant.  Once the slag/cement was off-loaded from the tanker delivery truck to the Site storage silo, the dust concentration returned to below the Response Limit.  Plastic sheeting will be placed over the delivery truck relief port with future deliveries to mitigate the dust.

If you have any questions or comments about the work on the Site, please use the Williamsburg MGP Telephone Hotline at (718) 403-3053.

Photos of Activity

Pre-trenching along the SOE wall alignment for the removal of shallow subsurface obstructions.

Pre-drilling for deep obstruction identification and removal along the SOE wall alignment.

Pre-drilling and pre-trenching for obstruction removal along the SOE wall alignment.

Loading lined trucks with impacted material for shipment to Clean Earth of North Jersey for treatment and disposal.

ISS BSM in the Holder No. 2 tank beneath the Temporary Containment Building (TCB).

Off-loading cement at the on Site grout plant silo for the ISS BSM.

Applying odor-suppressant foam to impacted material for odor control.

Pressure washing truck tires on the decontamination pad prior to exiting the Site onto North 11th Street.